OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 3y = 4 3x – 5y = –4

OpenStudy (anonymous):

Not sure what you meant by the elimination method. I would just add them together to obtain 8x-8y=0, hence x = y. Sub in x for y or vice versa for 5x - 3y = 4 or 3x - 5y = -4. x = y = 2

OpenStudy (anonymous):

What are the solutions?

OpenStudy (anonymous):

??????????

OpenStudy (anonymous):

x=2 y=2

OpenStudy (anonymous):

Okay, good I got it right.